17 September 2013: Tuesday

LDAP:

  • Authority:
    • Changed the permissions the COM/IS DN has to read entries as they coalesce their accounts with ours.
  • Issues:
    • There were no issues in last night’s Update! ¬†(OMG!!!)
  • 2.4.36:
    • Test continues, no issues with it on test primary and replica servers.
    • There have been reports of issues on the mailing list, so I’m waiting longer than normal before I push it into production.

Backups:

  • New Hardware:
    • Backups of the test subject ran at a calculated rate of 11.05 MB/s last night. ¬†That is compared to a range of 1.1 to 16.9 (Average of 8.77) over the 5 previous days to the old storage.
    • Loading up the other two new storage nodes.
  • Maintenance:
    • Planning next week’s maintenance
      • Working with Qualstar about firmware upgrade

Tags: ,

Comments are closed.